OpenStudy (anonymous):

Can someone help me with sinx(x) + sin(5x) =2 without using deductive method, please

OpenStudy (anonymous):

What is destructive method?? oops, sorry what is deductive method?

OpenStudy (owlfred):

I believe you mean \(\sin(x)+\sin(5x)=2\)? One way is to see that the sum can only be 2 if both \(\sin(x) \) and \(\sin(5x)\) are equal to 1.

OpenStudy (kainui):

Yeah, good point.